OpenStudy (anonymous):

Ball A is dropped from the top of a building of height h at the same instant that ball B is thrown vertically upward from the ground. When the balls collide, they are moving in opposite directions, and the speed of A is twice the speed of B. At what height does the collision occur? (Use any variable or symbol stated above as necessary.)

OpenStudy (anonymous):

so it goes like this... We denote the height of the building as h, the position of ball A as yA, the initial position of ball A as yA0, and analogously for ball B. The equation of motion for ball A is the following: yA - yA0 = vA0t - (0.5)gt2 yA0 = h vA0 = 0 g = 9.81 m/s2 Therefore, yA = -(0.5)(9.81 m/s2)t2 + h => yA = -4.905t2 + h The equation of motion for ball B is the following: yB - yB0 = vB0t - (0.5)gt2 yB0 = 0 g = 9.81 m/s2 Therefore, yB = vB0t - (0.5)(9.81 m/s2)t2 => yB = vB0t - 4.905t2 The two balls collide when yA = yB: yA = yB => -4.905t2 + h = vB0t - 4.905t2 => h = vB0t For ball A: vA = vA0 - gt vA0 = 0 Therefore, vA = - gt => vA = -9.81t For ball B: vB = vB0 - gt Therefore, vB = vB0 - gt => vB = vB0 - 9.81t We also know that when yA = yB, vA = -2vB: vA = -2vB => -9.81t = -2[vB0 - 9.81t] => -9.81t = -2vB0 + 19.62t => 2vB0 = 29.43t => vB0 = 14.715t Since h = vB0t, we have: h = (14.715t)t = 14.715t2 Unfortunately, I couldn't figure out what to do from this point until I spent the better part of the last hour typing this thread. Then, as I was typing the previous sentence, it hit me: xA = -4.905t2 + h = -4.905t2 + 14.715t2 => xA = 9.81t2 Therefore, xA / h = 9.81t2 / (14.715t2) => xA / h = 2/3 => xA = 2h/3, Hope it helps.